liquid digoxin is available as 0.05mg/ml. How much would you need for a dose of 36 mcg?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

        Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

   Here

        Desired dose(D) = 36 mcg.

        Dose in hand(H) = 0.05 mg

Convert mg to mcg for calculation

         1 mg = 1000 mcg

So

        Dose in hand (H) = 0.05 * 1000

                                     = 50 mcg.

               Quantity(Q) = 1 ml.

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 36/50 * 1

                                                      = 0.7 * 1

                                                      = 0.7 ml.